# Conformalized Credal Regions for Classification with Ambiguous Ground Truth

### v1

@@ Abstract.
An open question in Imprecise Probabilistic Machine Learning is how to empirically derive a credal region (i.e., a closed and convex family of probabilities on the output space) from the available data, without any prior knowledge or assumption. In classification problems, credal regions are a tool that is able to provide provable guarantees under realistic assumptions by characterizing the uncertainty about the distribution of the labels. Building on previous work, we show that credal regions can be directly constructed using conformal methods. This allows us to provide a novel extension of classical conformal prediction to problems with ambiguous ground truth, that is, when the exact labels for given inputs are not exactly known. The resulting construction enjoys desirable practical and theoretical properties: (i) conformal coverage guarantees, (ii) smaller prediction sets (compared to classical conformal prediction regions) and (iii) disentanglement of uncertainty sources (epistemic, aleatoric). We empirically verify our findings on both synthetic and real datasets.
1. Introduction
@@
In most real-world applications of machine learning, especially in the context of safety-critical applications such as healthcare, researchers have found it difficult to reason with precise probabilities. Instead, practitioners have become comfortable reasoning in terms of families of probabilities, often in the form of closed and convex sets called credal regions. Imprecise probabilistic machine learning (IPML) [22, 56, 23, 7, 9, 10, 24, 38, 5, 6, 11, 13, 46] aims to develop machine learning theory and methods that work with such imprecise probabilities. Such tools allow to better quantify and disentangle different types of uncertainty, e.g., epistemic (model) and aleatoric (data) uncertainties, which play key roles in any machine learning system. However, other sources of uncertainty are also highly relevant; for example, uncertainty originating from the annotation process used to derive ground truth labels [51].
@@
A crucial line of research in IPML is that of empirically deriving credal regions without any prior knowledge or assumption. The first steps in this direction were made recently by [14, 15]. They discovered that, subject to a so-called consonance assumption (see Section 4.2) and given exchangeable calibration data, the conformal transducer assigning a $p$-value to each possible label uniquely identifies a credal region. This is a very promising result, since this only relies on having to select a non-conformity measure, and the obtained coverage guarantee is valid irrespective of this choice. Unfortunately, [14, 15] do not provide an implementation of their results on credal regions in real-world, complex datasets. Moreover, they ignore the particularly interesting case of ambiguous ground truth [51, 52] where labels are not crisp, but subject to uncertainty due to rater disagreement [55, 57] and imperfect labeling tools. In fact, calibration data are often not exactly classified; we can see this in machine learning [18, 50], but especially in medicine [25, 41, 45, 48] and natural language processing [43].
@@
Contributions. In this paper, we follow an alternative way of constructing credal regions in a conformal way, inspired by recent work [31, 32, 35, 40]. Specifically, we directly construct conformalized families of probabilities and show that these are in fact credal regions, i.e. convex and closed. Furthermore, we extend this approach to classification problems with ambiguous ground truth, first studied in the context of conformal prediction by [52, 53]. [^1] Reference [52] refers to the published version, while [53] to the first version of the paper. Since many of the concept needed in the present paper were not included in the final, published version, we differentiate between the two for ease of reference. These problems are extremely relevant from an applied point of view. Instead of a “precise” calibration set $(\mathbf{x^{n}},\mathbf{y^{n}})=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}\subset\mathcal{X}\times\mathcal{Y}$, [^2] We use lower case letters for realizations, while capital letters for random variables. In this case, calibration data $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ are realizations that we observe, while the test example $X_{n+1}$ is a random variable that has not yet been realized. we consider a calibration set $D_{\text{c}}=\{(x_{i},\lambda_{i})\}$ that encompasses the ground truth ambiguity, where the $\lambda_{i}$’s are probability vectors, and the $k$-th entry $\lambda_{i,k}$ is the probability of $k$ being the correct label for $x_{i}$. [^3] A pair $(x_{i},\lambda_{i})$ in the calibration set is the realization of a random pair $(X_{i},\Lambda_{i})$. To see that this is a generalization of the usual classification problems, notice that we can write a pair $(x_{i},y_{i})$ as a pair $(x_{i},\lambda_{i})$, where $\lambda_{i}$ is a one-hot encoding vector with entry $1$ on the correct label $k=y_{i}$. By directly conformalizing in the probability space, we obtain an appealing calibration property for the constructed credal region: the true data generating process belongs to the credal region with high probability $1-\alpha$, where $\alpha$ is chosen by the user. Due to this property, our credal regions are well-suited to disentangle and quantify aleatoric and epistemic uncertainties associated with the analysis at hand [30, 47]. On both synthetic and real datasets, we verify both the coverage of the constructed regions and the true label coverage guarantee of the derived predictive sets. In contrast to [14, 15], we do not require any assumption beyond exchangeability. Using imprecise highest density sets, we can derive predictive sets of labels that are more efficient, i.e., smaller on average, compared to those by [53].
@@
Outline. This paper is structured as follows: First, in Section 2, we recall the plausibility regions and the conformal predictive sets derived in [53] and define the corresponding conformal coverage guarantee which we built on in the following. Then, in Section 3 we revisit the necessary background on imprecise probabilistic concepts relevant for this paper, including imprecise highest density sets (IHDSs). Part of our main results, Section 4 shows that the plausibility region in [53] is equivalent to a credal region, which also satisfies a desirable calibration property. Then, Sections 4.1 and 4.2 discuss the utility of our credal region the field of Imprecise Probabilistic Machine Learning in general and the relation and improvements over the work of [14, 15] in particular. Finally, Section 5 presents our findings on obtaining more efficient prediction sets from our credal regions compared to [53]. We present our experimental evidence in Section 6 and conclude in Section 7.
2. Conformal Plausibility Regions
@@
In [52, 53], the authors study conformal prediction for $K$-class classification in the context of ambiguous ground truth. Standard split conformal prediction typically assumes instead a calibration set of examples with “crisp” ground truth labels $\{(X_{i},Y_{i})\}_{i=1}^{n}$ [49]. Realizing that such crisp ground truth labels might not be available in many practical settings, [52, 53] assume a calibration set of examples and so-called plausibilities, $\{(X_{i},\Lambda_{i})\}_{i=1}^{n}$, where plausibilities $\lambda_{i}$ (i.e., the realizations of $\Lambda_{i}$) represent categorical distributions over the $K$ possible labels. [^4] They may be, for example, realizations from random variables $\Lambda_{i}$ distributed according to a Dirichlet distribution. This allows to represent ambiguous examples where the corresponding distribution $\mathbb{P}[Y|X]$ is not one-hot and might have high entropy. In the full-information setting, these plausibility vectors may represent the true distribution $\mathbb{P}[Y|X]$ directly; in many practical scenarios, however, we can only approximate the true distribution. For example, disagreement among annotators frequently indicates ambiguity and deterministic or probabilistic aggregation of multiple annotators can be used as plausibilities [51]. Then, we assume the true distribution can be obtained in the limit of infinite “faithful” annotators.
@@
where $\Delta^{K-1}$ is the unit simplex in $\mathbb{R}^{K}$, and $e(X_{n+1},\Lambda_{n+1}=\lambda)\coloneqq\sum_{k=1}^{K}\lambda_{k}E(X_{n+1},k)$ is a score derived from a so-called conformity score $E(X_{n+1},k)$ based on the probabilistic model prediction $p_{k}(X_{n+1})$. Such a model (e.g. the softmax output for class $k$ of a trained neural network) approximates the posterior class probabilities, $p_{k}(X_{n+1})\approx\mathbb{P}(Y_{n+1}=k\mid X_{n+1},D)$, where $D$ denotes the training set, and it is assumed to be available. Conformity score $E$ is built so that a high score is more unlikely, hence we can interpret $E$ as a function assigning a score to the assertion “$k$ is the correct label for the realization $x_{n+1}$ of test example $X_{n+1}$”, which is lower the more the pair $(x_{n+1},k)$ “lacks conformity” with the training data. It is worth noting that alternative definitions of $e$ have also recently been explored by [32]. The threshold $\tau$ is then chosen using a simple quantile computation on the calibration set,
@@
Under the assumption that $\{(X_{i},\Lambda_{i})\}_{i=1}^{n+1}$ are exchangeable, following standard conformal prediction literature [53, Equation (13)], the plausibility region $c(X_{n+1})$ provides a coverage guarantee stating that $\mathbb{P}[\Lambda_{n+1}\in c(X_{n+1})]\geq 1-\alpha$, [^5] This probability is also on the calibration set, since the latter is used to derive $c(X_{n+1})$. where $\Lambda_{n+1}$ is the unobserved true plausibility vector of the test example. If the $\lambda_{i}$’s in the calibration set correspond to the true distributions $\mathbb{P}[Y|X_{i}=x_{i}]$, this provides coverage with respect to the true distribution. As [53, Section 3.1] details, however, in many practical settings, the plausibilities are obtained from expert annotations. With finite annotators, the authors are able to give a coverage guarantee, $\mathbb{P}_{\text{agg}}[\Lambda_{n+1}\in c(X_{n+1})]\geq 1-\alpha$, stated in terms of the distribution $\mathbb{P}_{\text{agg}}$ that explicitly captures how annotations are aggregated into plausibilities. The underlying assumption is that $\lambda_{n+1,k}=\mathbb{P}_{\text{agg}}[Y_{n+1}=k|X_{n+1}]$ where, ideally, $\mathbb{P}_{\text{agg}}\approx\mathbb{P}$, and the annotators are the same for the calibration set and the test example. In any case, annotators can “agree to disagree” such that $\lambda_{n+1}$ has high entropy in which case $x_{n+1}$ is called ambiguous. For simplicity, we ignore this caveat for the presentation of this paper and write $\mathbb{P}$ henceforth.
The plausibility region $c(X_{n+1})$ is then used in [53, Equation (41)] to derive “plausibility-reduced” predictive sets (PRPS)
@@
Here $\lambda^{\sigma}=(\lambda_{\sigma_{1}},\ldots,\lambda_{\sigma_{K}})^{\top}$ corresponds to vector $\lambda$ sorted in descending order. In [53], this is also contrasted with regular conformal predictive sets (CPS) that are obtained by calibrating a threshold $\kappa$ on the per-label scores $E$ and constructing the CPS $C(X_{n+1})$ as $\{k\in\mathcal{Y}:E(X_{n+1},k)\geq\kappa\}$. There, the latter is further adapted to allow for ambiguous examples with plausibilities $\lambda_{i}$ available for calibration, see [53, Algorithm 1]. These adapted CPS $C(X_{n+1})$ are shown to be generally more efficient, i.e., smaller, compared to $\Psi(c(X_{n+1}))$ but unable to capture or even disentangle different sources of uncertainty. We make a step towards addressing this gap by deriving narrower predictive sets that allow for uncertainty quantification and disaggregation, using methods from imprecise probability.
3. Imprecise Probability
Following [14, 15] and additional previous work on IPML [3, 12, 17], we briefly introduce the notions of lower and upper probabilities, and also the concept of Imprecise Highest Density Sets (IHDS). These will play a pivotal role in the paper – and especially in Section 5 – as a way to derive predictive sets of labels from our credal regions, that are shown to be more efficient than the predictive sets $\Psi(c(X_{n+1}))$.
@@ 3. Imprecise Probability
Definition 1 (Imprecise Highest Density Set, [17]).
@@
In order to compute IHDSs in practice, we need another important result – proved in [3, Section 4.4] and [19].
@@
Since $\underline{P}$ is the lower envelope of a credal region, the sure loss avoidance condition is always met, as proven in [54]. We make this explicit in the following Lemma.
@@
Definition 1 can be used to derive a parallel between (generic) Conformal Prediction Sets (CPSs) and IHDSs. Condition (i) tells us that $Y_{n+1}$ belongs to $\text{IS}_{\mathcal{P},\delta}$ with $P$-probability of at least $1-\delta$, for all distributions $P$ in the credal region $\mathcal{P}$. Condition (ii) tells us that $\text{IS}_{\mathcal{P},\delta}$ is “efficient”. That is, it is the narrowest subset of the label set $\mathcal{Y}$ that is able to ensure the probabilistic guarantee of condition (i). Compared to CPSs, this is a weaker guarantee: First and foremost, the conformal guarantee is uniform, holding for all possible exchangeable distributions $P$ on $\mathcal{Y}$. Second, without assumptions on the credal region $\mathcal{P}$, condition (i) ignores whether the true distribution belongs to the credal set $\mathcal{P}$.
4. Conformal Plausibility Regions as Credal Regions
A key contribution of this paper is relating the plausibility regions $c(X_{n+1})$ in (1) to the imprecise probabilistic notion of credal regions. As we will show, this leads to a remarkable synergy that allows us to construct a credal regions for $X_{n+1}$ in a conformal way, and subsequently to use IHDSs to construct predictive sets of labels. Compared to [14, 15], our credal regions provide coverage, only requiring the exchangeability of $\{(X_{i},\Lambda_{i})\}_{i=1}^{n+1}$. In addition, we improve over [53]: the predictive sets that we derive are more efficient, i.e., always non-broader and sometimes strictly narrower. For a start, we show that the plausibility regions $c(X_{n+1})$ are convex and closed, and thus proper credal regions.

### v2

@@ Abstract
An open question in Imprecise Probabilistic Machine Learning is how to empirically derive a credal region (i.e., a closed and convex family of probability measures on the output space) from the available data, without any prior knowledge or assumption. In classification problems, credal regions are a tool that is able to provide provable guarantees under realistic assumptions by characterizing the uncertainty about the distribution of the labels. Building on previous work, we show that credal regions can be directly constructed using conformal methods. This allows us to provide a novel extension of classical conformal prediction to problems with ambiguous ground truth, that is, when the exact labels for given inputs are not exactly known. The resulting construction enjoys desirable practical and theoretical properties: (i) conformal coverage guarantees, (ii) smaller prediction sets (compared to classical conformal prediction regions) and (iii) disentanglement of uncertainty sources (epistemic, aleatoric). We empirically verify our findings on both synthetic and real datasets.
1 Introduction
@@
In most real-world applications of machine learning, especially in the context of safety-critical applications such as healthcare, researchers have found it difficult to reason with precise probabilities. Instead, practitioners have become comfortable reasoning in terms of families of probabilities, often in the form of closed and convex sets called credal regions. Imprecise probabilistic machine learning (IPML) (Denoeux, 2000; Zaffalon, 2002; Destercke et al., 2008; Caprio & Gong, 2023; Caprio & Mukherjee, 2023a; b; Dutta et al., 2023; Lu et al., 2024; Caprio, 2024; Caprio et al., 2024a; c; d; Sale et al., 2024) aims to develop machine learning theory and methods that work with such imprecise probabilities. Such tools allow to better quantify and disentangle different types of uncertainty, e.g., epistemic (model) and aleatoric (data) uncertainties, which play key roles in any machine learning system. However, other sources of uncertainty are also highly relevant; for example, uncertainty originating from the annotation process used to derive ground truth labels (Stutz et al., 2023a).
@@
A crucial line of research in IPML is that of empirically deriving credal regions without any prior knowledge or assumption. The first steps in this direction were made recently by Cella & Martin (2022a; b). They discovered that, subject to a so-called consonance assumption (see Section 3.2) and given exchangeable calibration data, the conformal transducer assigning a $p$-value to each possible label uniquely identifies a credal region. This is a very promising result, since this only relies on having to select a non-conformity measure, and the obtained coverage guarantee is valid irrespective of this choice. Unfortunately, Cella & Martin (2022a; b) do not provide an implementation of their results on credal regions in real-world, complex datasets. Moreover, they ignore the particularly interesting case of ambiguous ground truth (Stutz et al., 2023a; b) where labels are not crisp, but subject to uncertainty due to rater disagreement (Yan et al., 2014; Zheng et al., 2017) and imperfect labeling tools. In fact, calibration data are often not exactly classified; we can see this in machine learning (Dawid & Skene, 1979; Smyth et al., 1994), but especially in medicine (Feinstein & Cicchetti, 1990; McHugh, 2012; Raghu et al., 2019; Schaekermann, 2020), natural language processing (Pavlick & Kwiatkowski, 2019; Nie et al., 2020; Baan et al., 2022; Plank, 2022), and computer vision (Peterson et al., 2019a).
@@
Contributions. In this paper, we follow an alternative way of constructing credal regions in a conformal way, inspired by recent work (Javanmardi et al., 2023; 2024; Lienen et al., 2023; Martin Bordini et al., 2023). Specifically, we directly construct conformalized families of probabilities and show that these are in fact credal regions, i.e. convex and closed. Furthermore, we extend this approach to classification problems with ambiguous ground truth, first studied in the context of conformal prediction by Stutz et al. (2023b; c). [^1] Reference Stutz et al. (2023b) refers to the published version, while Stutz et al. (2023c) to the first version of the paper. Since many of the concept needed in the present paper were not included in the final, published version, we differentiate between the two for ease of reference. These problems are extremely relevant from an applied point of view. Instead of a “precise” calibration set $(\mathbf{x^{n}},\mathbf{y^{n}})=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}\subset\mathcal{X}\times\mathcal{Y}$, [^2] We use lower case letters for realizations, while capital letters for random variables. In this case, calibration data $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ are realizations that we observe, while the test example $X_{n+1}$ is a random variable that has not yet been realized. we consider a calibration set $D_{\text{c}}=\{(x_{i},\lambda_{i})\}$ that encompasses the ground truth ambiguity, where the $\lambda_{i}$’s are probability vectors, and the $k$-th entry $\lambda_{i,k}$ is the probability of $k$ being the correct label for $x_{i}$. [^3] A pair $(x_{i},\lambda_{i})$ in the calibration set is the realization of a random pair $(X_{i},\Lambda_{i})$. To see that this is a generalization of the usual classification problems, notice that we can write a pair $(x_{i},y_{i})$ as a pair $(x_{i},\lambda_{i})$, where $\lambda_{i}$ is a one-hot encoding vector with entry $1$ on the correct label $k=y_{i}$. By directly conformalizing in the probability space, we obtain an appealing calibration property for the constructed credal region: the true data generating process belongs to the credal region with high probability $1-\alpha$, where $\alpha$ is chosen by the user. Due to this property, our credal regions can be used to disentangle and quantify aleatoric and epistemic uncertainties associated with the analysis at hand (Hüllermeier & Waegeman, 2021; Sale et al., 2023). On both synthetic and real datasets, we verify both the coverage of the constructed regions and the true label coverage guarantee of the derived predictive sets. In contrast to Cella & Martin (2022a; b), we do not require any assumption beyond exchangeability. Using imprecise highest density sets, we can derive predictive sets of labels that are more efficient, i.e., smaller on average, compared to those by Stutz et al. (2023c).
@@
Outline. This paper is structured as follows: Section 2 provides the required background. First, in Section 2.1, we recall the plausibility regions and the conformal predictive sets derived in Stutz et al. (2023c) and define the corresponding conformal coverage guarantee which we built on in the following. Then, in Section 2.2 we revisit the necessary imprecise probabilistic concepts relevant for this paper, including imprecise highest density sets (IHDSs). Part of our main results, Section 3 shows that the plausibility region in Stutz et al. (2023c) is equivalent to a credal region, which also satisfies a desirable calibration property. Then, Sections 3.1 and 3.2 discuss the utility of our credal region the field of Imprecise Probabilistic Machine Learning in general and the relation and improvements over the work of Cella & Martin (2022a; b) in particular. Finally, Section 4 presents our findings on obtaining more efficient prediction sets from our credal regions compared to (Stutz et al., 2023c). We present our experimental evidence in Section 5 and conclude in Section 6.
2 Background
2.1 Conformal Plausibility Regions
@@
In Stutz et al. (2023b; c), the authors study conformal prediction for $K$-class classification in the context of ambiguous ground truth. Standard split conformal prediction typically assumes instead a calibration set of examples with “crisp” ground truth labels $\{(X_{i},Y_{i})\}_{i=1}^{n}$ (Shafer & Vovk, 2008). Realizing that such crisp ground truth labels might not be available in many practical settings, Stutz et al. (2023b; c) assume a calibration set of examples and so-called plausibilities, $\{(X_{i},\Lambda_{i})\}_{i=1}^{n}$, where plausibilities $\lambda_{i}$ (i.e., the realizations of $\Lambda_{i}$) represent categorical distributions over the $K$ possible labels. [^4] They may be, for example, realizations from random variables $\Lambda_{i}$ distributed according to a Dirichlet distribution. This allows us to represent ambiguous examples where the corresponding distribution $\mathbb{P}[Y|X]$ is not one-hot and might have high entropy. In the full-information setting, these plausibility vectors may represent the true distribution $\mathbb{P}[Y|X]$ directly; in many practical scenarios, however, we can only approximate the true distribution. For example, disagreement among annotators frequently indicates ambiguity and deterministic or probabilistic aggregation of multiple annotators can be used as plausibilities (Stutz et al., 2023a). Then, we assume the true distribution can be obtained in the limit of infinite “faithful” annotators.
@@
where $\Delta^{K-1}$ is the unit simplex in $\mathbb{R}^{K}$, and $e(X_{n+1},\Lambda_{n+1}=\lambda)\coloneqq\sum_{k=1}^{K}\lambda_{k}E(X_{n+1},k)$ is a score derived from a so-called conformity score $E(X_{n+1},k)$ based on the probabilistic model prediction $p_{k}(X_{n+1})$. Such a model (e.g. the softmax output for class $k$ of a trained neural network) approximates the posterior class probabilities, $p_{k}(X_{n+1})\approx\mathbb{P}(Y_{n+1}=k\mid X_{n+1},D)$, where $D$ denotes the training set, and it is assumed to be available. Conformity score $E$ is built so that a high score is more unlikely, hence we can interpret $E$ as a function assigning a score to the assertion “$k$ is the correct label for the realization $x_{n+1}$ of test example $X_{n+1}$”, which is lower the more the pair $(x_{n+1},k)$ “lacks conformity” with the training data. It is worth noting that alternative definitions of $e$ have also recently been explored by Javanmardi et al. (2024). The threshold $\tau$ is then chosen using a simple quantile computation on the calibration set,
@@
Under the assumption that $\{(X_{i},\Lambda_{i})\}_{i=1}^{n+1}$ are exchangeable, following standard conformal prediction literature (Stutz et al., 2023c, Equation (13)), the plausibility region $c(X_{n+1})$ provides a coverage guarantee stating that $\mathbb{P}[\Lambda_{n+1}\in c(X_{n+1})]\geq 1-\alpha$, [^5] This probability is also on the calibration set, since the latter is used to derive $c(X_{n+1})$. where $\Lambda_{n+1}$ is the unobserved true plausibility vector of the test example. If the $\lambda_{i}$’s in the calibration set correspond to the true distributions $\mathbb{P}[Y|X_{i}=x_{i}]$, this provides coverage with respect to the true distribution. As (Stutz et al., 2023c, Section 3.1) details, however, in many practical settings, the plausibilities are obtained from expert annotations. With finite annotators, the authors are able to give a coverage guarantee, $\mathbb{P}_{\text{agg}}[\Lambda_{n+1}\in c(X_{n+1})]\geq 1-\alpha$, stated in terms of the distribution $\mathbb{P}_{\text{agg}}$ that explicitly captures how annotations are aggregated into plausibilities. The underlying assumption is that $\lambda_{n+1,k}=\mathbb{P}_{\text{agg}}[Y_{n+1}=k|X_{n+1}]$ where, ideally, $\mathbb{P}_{\text{agg}}\approx\mathbb{P}$, and the annotators are the same for the calibration set and the test example. In any case, annotators can “agree to disagree” such that $\lambda_{n+1}$ has high entropy in which case $x_{n+1}$ is called ambiguous. For simplicity, we ignore this caveat for the presentation of this paper and write $\mathbb{P}$ henceforth. The interested reader can find some extra details on the difference between $\mathbb{P}_{\text{agg}}$ and $\mathbb{P}$ in Appendix A.
The plausibility region $c(X_{n+1})$ is then used in (Stutz et al., 2023c, Equation (41)) to derive “plausibility-reduced” predictive sets (PRPS) for some user-chosen $\delta\in[0,1]$,
@@
Here $\lambda^{\sigma}=(\lambda_{\sigma_{1}},\ldots,\lambda_{\sigma_{K}})^{\top}$ corresponds to vector $\lambda$ sorted in descending order. Equation 2 tells us that $\Psi(c(X_{n+1}))$ is the set of all labels $y$ that are assigned a probability non-smaller than $1-\delta$ (of being the correct one for the realization of $X_{n+1}$), by at least one probability vector $\lambda$ that belongs to the plausibility region $c(X_{n+1})$.
In Stutz et al. (2023c), this is also contrasted with regular conformal predictive sets (CPS) that are obtained by calibrating a threshold $\kappa$ on the per-label scores $E$ and constructing the CPS $C(X_{n+1})$ as $\{k\in\mathcal{Y}:E(X_{n+1},k)\geq\kappa\}$. There, the latter is further adapted to allow for ambiguous examples with plausibilities $\lambda_{i}$ available for calibration, see (Stutz et al., 2023c, Algorithm 1). These adapted CPS $C(X_{n+1})$ are shown to be generally more efficient, i.e., smaller, compared to $\Psi(c(X_{n+1}))$, but are unable to capture or even disentangle different sources of uncertainty. We make a step towards addressing this gap by deriving narrower predictive sets that allow for uncertainty quantification and disaggregation, using methods from imprecise probability.
2.2 Imprecise Probability
Following Cella & Martin (2022a; b) and additional previous work on IPML (Augustin et al., 2014; Caprio & Seidenfeld, 2023; Coolen, 1992), we briefly introduce the notions of lower and upper probabilities, and also the concept of Imprecise Highest Density Sets (IHDS). These will play a pivotal role in the paper – and especially in Section 4 – as a way to derive predictive sets of labels from our credal regions, that are shown to be more efficient than the predictive sets $\Psi(c(X_{n+1}))$.
@@ 2.2 Imprecise Probability
Definition 1 (Imprecise Highest Density Set, Coolen (1992)).
@@
In order to compute IHDSs in practice, we need another important result – proved in (Augustin et al., 2014, Section 4.4) and de Campos et al. (1994).
@@
Since $\underline{P}$ is the lower envelope of a credal region, the sure loss avoidance condition is always met, as proven in Walley (1991). We make this explicit in the following Lemma.
@@
Definition 1 can be used to derive a parallel between (generic) Conformal Prediction Sets (CPSs) and IHDSs. Condition (i) tells us that $Y_{n+1}$ belongs to $\text{IS}_{\mathcal{P},\delta}$ with $P$-probability of at least $1-\delta$, for all distributions $P$ in the credal region $\mathcal{P}$. Condition (ii) tells us that $\text{IS}_{\mathcal{P},\delta}$ is “efficient”. That is, it is the narrowest subset of the label set $\mathcal{Y}$ that is able to ensure the probabilistic guarantee of condition (i). Compared to CPSs, this is a weaker guarantee: First and foremost, the conformal guarantee is uniform, holding for all possible exchangeable distributions $P$ on $\mathcal{Y}$. Second, without assumptions on the credal region $\mathcal{P}$, condition (i) ignores whether the true distribution belongs to the credal region $\mathcal{P}$.
3 Conformal Plausibility Regions as Credal Regions
A key contribution of this paper is relating the plausibility regions $c(X_{n+1})$ in equation 1 to the imprecise probabilistic notion of credal regions. As we will show, this leads to a remarkable synergy that allows us to construct a credal regions for $X_{n+1}$ in a conformal way, and subsequently to use IHDSs to construct predictive sets of labels. Compared to Cella & Martin (2022a; b), our credal regions provide coverage, only requiring the exchangeability of $\{(X_{i},\Lambda_{i})\}_{i=1}^{n+1}$. In addition, we improve over Stutz et al. (2023c): the predictive sets that we derive are more efficient, i.e., always non-broader and sometimes strictly narrower. For a start, we show that the plausibility regions $c(X_{n+1})$ are convex and closed, and thus proper credal regions.